OpenStudy (anonymous):

A polyhedron has 10 faces and 24 edges. What is the number of vertices of the polyhedron?

OpenStudy (anonymous):

12 16 10 14

OpenStudy (amistre64):

there is a famous formula for this

OpenStudy (amistre64):

i can never remember it tho :) something like:\[v+f=2-e\]

OpenStudy (amistre64):

v+f = 2+e maybe? that sounds more righter

Parth (parthkohli):

\[F + V = E + 2 \]This is commonly known as Euler's Formula. What I stated above may also be written as\[ F + V - E = 2\]

OpenStudy (anonymous):

Thank you!